The admission fee at an amusement park is $1.50 for children and $4 for adults. On a certain day, 236 people entered the park, and the admission fees collected totaled 604.00 dollars. How many children and how many adults were admitted?

Your answer is
number of children equals
number of adults equals

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

x children
236-x adults
money equation 1.50x+4(236-x)=604
1.50x+944-4x=604
2.5x=340
x=136 children ($204
236-x=100 adults ($400)

Answer by greenestamps(13200)



The other tutor showed a typical formal algebraic solution method.

If formal algebra is not required and the speed of getting the answer is important, you can try this informal solution method using logical reasoning and simple mental arithmetic.

(1) If all 236 tickets were child tickets, the total sales would have been 236(1.50) = 354.00
(2) The actual total, 604.00, is 604-354 = 250.00 more
(3) The difference in cost between a child ticket and an adult ticket is 4.00-1.50 = 2.50
(4) The number of adult tickets needed to make the additional 250.00 is 250.00/2.50 = 100

ANSWER: 100 adults and 236-100 = 136 child tickets

CHECK: 100(4)+136(1.5)=400+204=604